# Autonomous Learning with High-Dimensional Computing Architecture Similar to von Neumann’s

### v1

@@
ABSTRACT: We model human and animal learning by computing with high-dimensional vectors ($H$ = 10,000 for example). The architecture resembles traditional (von Neumann) computing with numbers, but the instructions refer to vectors and operate on them in superposition. The architecture includes a high-capacity memory for vectors, analogue of the random-access memory (RAM) for numbers. The model’s ability to learn from data reminds us of deep learning, but with an architecture closer to biology. The architecture agrees with an idea from psychology that human memory and learning involve a short-term working memory and a long-term data store. Neuroscience provides us with a model of the long-term memory, namely, the cortex of the cerebellum. With roots in psychology, biology, and traditional computing, a theory of computing with vectors can help us understand how brains compute. Application to learning by robots seems inevitable, but there is likely to be more, including language. Ultimately we want to compute with no more material and energy than used by brains. To that end, we need a mathematical theory that agrees with psychology and biology, and is suitable for nanotechnology. We also need to exercise the theory in large-scale experiments. Computing with vectors is described here in terms familiar to us from traditional computing with numbers.
KEYWORDS: machine learning; holographic reduced representation; hyperdimensional computing; vector-symbolic architecture; associative memory; Marr–Albus model; cerebellum; sparse distributed memory; working memory
@@ 1 Background
Neural-net associative memories for high-dimensional vectors were studied already in the 1970s, but operations of the right kind for computing with vectors were found only in the 1990s (Plate 1991 for real and complex vectors, Kanerva 1996 for binary vectors, and Gayler & Wales 1998 for vectors of integers). Computing with vectors is subtle in that it takes place in superposition. We compare computing with vectors to conventional computing with numbers and argue for modeling and simulating human and animal cognition with a von Neumann-like architecture for computing with high-dimensional vectors.
We begin with an overview of conventional computing and identify corresponding circuits for computing with vectors. We then discuss reasons to distinguish between a working memory and a long-term memory. The vector math is carried out in the working memory and the results are stored in the long-term memory and ultimately drive action. We make no attempt to map the working memory to neural circuits in the brain, but observe that the circuits in the cerebellum are laid out beautifully in 3D for a massive long-term memory for high-dimensional vectors.
The rest of the paper and the target article are about applying the architecture to autonomous learning by robots, modeled after learning by humans and animals with sensory–motor systems controlled by advanced brains. The paper represents a systems approach that is aligned with James Albus’s ideas about robots with human-like faculties. His book on Brains, Behavior, and Robotics (1981) was published a decade before ideas about computing with vectors in superposition began to take hold, but is still most relevant and illuminating.
@@ 2 Traditional Computing with Numbers
This section is meant to help readers who know computers from having programmed them in high-level languages but know less about the hardware within. The lesson is that the computer in itself is exceedingly simple and that its success comes from programming.
The traditional computing architecture, known also as the von Neumann architecture, consists of a memory for storing programs and data (Random-Access Memory, RAM), and a processor that runs the program and does the math (Central Processing Unit, CPU). The memory is made of sequentially numbered locations, and a location’s number serves as its address. Even small computers have memories with millions of locations.
A memory location is a piece of hardware that can hold a small amount of data, for example a number, a letter, or a simple command for the processor to carry out, such as adding two numbers. A memory location can be visualized as a fixed number of bits, 32 for example, and the programmer decides what the bits in any particular location mean.
The CPU has two functions, step-by-step running of the program and the math that is built into the computer’s circuits. The math usually includes addition, subtraction, multiplication, and division of numbers, and logic operations on strings of bits (AND, OR, NOT, Shift and Rotate). Fancier operations are made with programs that combine these basic operations. For example, a program to calculate the mean of a set of numbers copies the numbers, one at a time, from the memory into the processor, adds it to the sum, divides the final sum by the count, and stores the result in memory for later use—for comparing the mean heights of two groups of people, for example.
Random Access is an all-important property of the memory. It means that the contents of any location are immediately available to the processor. The processor merely needs to “know” the location’s address. Random access is in contrast to storing data on tape or disk that make the processor wait for the addressed memory location to be under the read/write head. The “miracle” of random access is achieved with an address-decoder circuit.
Random access allows lists of numbers to be stored so that any number in the list can start the retrieval of the rest of the list (if the numbers are addresses of memory locations). The numbers in the list are used to address the memory and as data to be stored, and the list is stored by stepping through it one number at a time: each number is then used as the address, to store the next number in the list. Linked lists, also called pointer chains, are standard fare in list-processing languages such as Lisp, and a useful way to store a system’s history, as discussed below.
@@ 3 Computing with Vectors
The traditional (von Neumann) architecture for computing with numbers is readily adapted to computing with high-dimensional fixed-width vectors (e.g., $H$ = 10,000). The new architecture includes an instruction set for vectors, a Hyperdimensional Processing Unit (HPU) to carry out the vector operations, and a random-access memory for vectors.
@@ 3.1 Instruction Set for Vectors
Three simple operations make a surprisingly powerful system to compute with: vector addition, coordinatewise multiplication, and the shuffling of vector coordinates (permutation). The vectors can be binary or integer or real or complex—the computing power comes more from high dimensionality (e.g., $H$ = 10,000) than from the nature of vector components.
The operations have been discussed at length in the literature (Plate 2003, Kanerva 2009, 2023) and are illustrated here with binary vectors. Addition is by coordinatewise majority rule with ties broken at random (equivalently, ordinary vector addition followed by applying a threshold to each coordinate); multiplication is coordinatewise Exclusive-Or (XOR); and permutations reorder vector coordinates. The algebra of addition and multiplication resemble ordinary arithmetic in regard to associativity, commutativity, invertibility, and distributivity. Permutations have no counterpart in ordinary arithmetic. They are invertible, most do not commute, and they distribute over both vector addition and multiplication. Permutations of coordinates add all finite groups up to size $H$ into the vector math.
The operations allow traditional data structures (sets, sequences, lists, trees, stacks, etc.) to be represented and operated on in superposed vectors (Kleyko et al. 2022). For example, sets can be encoded with addition, the binding of variables to values with multiplication, sequences with permutations, and data structures at large with combinations of the three operations. For a simple example, a vector that means $x=a$ can be encoded by the product $P=X*A$ where $X$ and $A$ are $H$-dimensional vectors representing the variable $x$ and the value $a$, and * is coordinatewise multiplication (XOR for binary vectors). The three operations are also used to decode a composed vector into its constituent vectors: the value of $x$ in $P$ can be recovered by multiplying $P$ with (the inverse of) $X$, namely $X*P=X*(X*A)=(X*X)*A=A$, and by noting that XOR is its own inverse so that $X*X$ evaluates to a vector of 0s. In ordinary computing, the variable $x$ is represented by an address of a memory location and is buried in the program code, and the value $a$ is represented by the contents of the location. The three operations also allow regularities in data to be captured in vectors, opening a path to statistical learning.
@@ 3.2 Computing in Superposition
Computing in superposition goes by many names, depending mainly on the intended audience—whether mathematicians, physicist, cognitive scientists, linguists, engineers, or computer scientists: Holographic Reduced Representation (HRR), Vector Symbolic Architecture (VSA), Context-Dependent Thinning, Hyperdimensional Computing, and Semantic Pointer Architecture (SPA). Each item or element or feature or thing or piece of data—each vector—that is included in a superposed vector is distributed evenly over all vector components—any subset of vector components represents the same thing as the entire vector, only less precisely. Representation of this kind is called “holographic” or “holistic,” as opposed to the conventional localist representation where a variable is represented by an address of a memory location and its value by the contents of the location. The encoding of $x=a$ with the vector $P$ above is an example of holistic representation.
@@
Computing in superposition begins with $H$-dimensional seed vectors chosen at random (with i.i.d. components), to represent basic entities such as variables and names. In essence, they are symbols to compute with. They can be used as addresses to memory and as data stored in memory, and as inputs to the math operations that combine them into new vectors, which in turn can address the memory, be stored as data, and serve as inputs to further vector operations. New representations are made from existing ones with explicit calculation, which is fundamentally different from the generation of representations in an autoencoder or in the layers of a deep neural net as it is trained.
@@
The operations allow a high degree of parallelism. Because addition and multiplication happen coordinatewise, independently of other coordinates, all $H$ (10,000) can be done at the same time. Furthermore, because the operations are simple, it is feasible for each dimension to have its own processor circuit. Computing with vectors in superposition can then be seen as each coordinate having its own simple computer, all computing the same thing but at exceedingly low precision—down to 1-bit. On the aggregate, however, they compute with high precision if the dimensionality is high enough—and even if some of the simple circuits malfunction! In contrast, traditional circuits for computing with numbers are complicated and are expected to work flawlessly. The simplicity of the coordinatewise operations makes it possible to distribute them throughout the system, reducing the need to move data before it can be computed on. This is called “in-memory computing.”
4 Ideas for Architecture
from Biology and Psychology
@@ 4 Ideas for Architecture
How does computing with vectors fare in the light of biology and behavioral sciences? Psychological research in the 1960s and ’70s identified two levels of organization of human memory, short-term working memory and long-term data store (Atkinson & Shiffrin 1968, Anderson & Bower 1973; see also Wikipedia article on “Atkinson–Shiffrin memory model”). Representations are assembled and analyzed in the working memory and stored in the long-term memory (Tulving & Donaldson 1972, Tulving & Thomson 1973). These early models summarize behavior without accounting for underlying mechanisms and their math.
@@
The psychologists’ models are similar to traditional computing that assembles and interprets data structures in the central processing unit (CPU) and stores them in the random-access memory (RAM). Computing with vectors follows the same organization: vectors are encoded and decoded in the working memory (analogue of CPU) and stored indefinitely in the long-term memory (analogue of RAM). We will discuss the long-term memory first.
@@ 4.1 Long-Term Memory for Vectors
Memory allows past experience to bear on the present situation. Associative memory, also called content-addressable (these terms lack a precise definition), has been an object of psychological inquiry for over a century (James 1890, Raaijmakers & Shiffrin 1981), and mathematical models have been proposed since the early days of artificial neural nets (see Hinton & Anderson 1981 for a review). Sparse Distributed Memory (Kanerva 1988b) takes the idea a step further by assuming that the memory could have billions of locations and that its function is long-term data storage. In the rest of this paper, “long-term memory,” “associative memory,” and “high-dimensional RAM” mean the same thing and are often referred to simply as the “memory.”
A psychologically realistic memory for vectors is also addressed by vectors—think of a 10,000-bits wide memory, addressed by 10,000-bit vectors. The number of possible addresses (2-to-power-10,000) is beyond astronomical; only a minuscule fraction can be present as physical storage locations. Yet it is possible to simulate random access by distributing data over multiple locations and reconstructing it upon retrieval. The idea is familiar from neural-net associative memories at large (e.g., Hopfield 1982).
@@
When the sparse distributed memory is addressed, a small number of locations are activated—fewer than one in a thousand. The activated locations either accept (store) data or retrieve (recall) data. There is a further difference to note: storing a number in a computer memory replaces the old contents of the activated/addressed location; storing a vector in an associative memory adds it to the contents of all activated locations—the newly stored data are superposed.
@@
An associative memory’s likeness to the cortex of the cerebellum is remarkable (Marr 1969, Albus 1971, Kanerva 1993). With a granule cell of the cerebellum representing a storage location, the human brain would have 50 billion locations in which to store a vector. A rough calculation shows that an associative memory of that size could store one vector every minute (more than 10 million bits a day) over a life span of a 100 years.
@@
Catastrophic forgetting is a known phenomenon of artificial neural nets. It is in part due to the limit on the amount of information that any given “hardware” can store, and in part to training a monolithic network with error back-propagation. Catastrophic means that when the capacity limit is reached, the net changes drastically and begins to produce outputs that are unrelated to the vectors it was trained with—it’s akin to a phase transition in physics where the properties of a substance change suddenly, as when super-cooled water turns into a block of ice. The alternative is graceful degradation, which can be achieved with a memory that is organized in locations whose contents can change but addresses are fixed, as they do in ordinary computer memory, in the sparse distributed memory, and apparently also in the cortex of the cerebellum where most synapses are modifiable (store data) but a small fraction (1%) is fixed (they decode the address). However, that 1% serves as a framework for recording a lifetime of experience.

### v2

@@
Abstract
We model human and animal learning by computing with high-dimensional vectors ($D$ = 10,000 for example). The architecture resembles traditional (von Neumann) computing with numbers, but the instructions refer to vectors and operate on them in superposition. The architecture includes a high-capacity memory for vectors, counterpart of the random-access memory (RAM) for numbers. The model’s ability to learn from data reminds us of deep learning, but with an architecture closer to biology. The architecture agrees with an idea from psychology that human memory and learning involve a short-term working memory and a long-term data store. Neuroscience provides us with a model of the long-term memory, namely, the cortex of the cerebellum. With roots in psychology, biology, and traditional computing, a theory of computing with vectors can help us understand how brains compute. Application to learning by robots seems inevitable, but there is likely to be more, including language. Ultimately we want to compute with no more material and energy than used by brains. To that end, we need a mathematical theory that agrees with psychology and biology, and is suitable for nanotechnology. We also need to exercise the theory in large-scale experiments. The analogy with traditional computing suggests that the architecture be programmable in terms of variables, values and data structures, the very things that have made traditional computing ubiquitous and that seem worth learning from and emulating.
1 Preamble
It seems best to clear a likely misconception up front. We generally agree that the everyday computer is a “von Neumann machine,” and we call other architectures “non-von-Neumann,” including Artificial Neural Nets (ANNs). Computing with wide vectors, as discussed in this paper, grew out of artificial neural nets and is an ANN of a kind, but it realizes a von Neumann-like architecture with an instruction set and a memory for wide vectors, and it allows variables and values to be used in traditional ways. Three vector operations are used to encode and decode composite structure, and to capture probabilities in bit counts and vector similarity. Minimal reliance on backpropagation is another thing that sets the “new” architecture apart from mainstream ANNs. In abstract theory terms, the architecture is no less powerful than an ordinary computer, or the Universal Turing Machine without an endless tape. The finite automaton at the core is simple and easily realized in an algebra of vectors. The point is to program with vectors as basic objects, much as we now program with numbers.
@@ 2 Background
Neural-net associative memories for high-dimensional (aka “wide”) vectors were studied already in the 1970s, but operations of the right kind for computing with vectors were found only in the 1990s, starting with Plate’s Holographic Reduced Representation (Plate 1991). Computing with vectors is subtle in that it takes place in superposition.
We compare computing with vectors to conventional computing with numbers and argue for modeling and simulating human and animal cognition with a von Neumann-like architecture for computing with vectors. We begin with an overview of conventional computing and identify corresponding circuits for computing with vectors. We then discuss reasons to distinguish between working memory and long-term memory. The vector math is carried out in the working memory and the results are stored in the long-term memory and ultimately drive action. We make no attempt to map the working memory to neural circuits in the brain, but observe that the circuits of the cerebellum are laid out beautifully in 3D for a massive long-term memory for wide vectors (see e.g. Eccles et al. (1967) and figures 7.2 and 7.3 of Llinás et al. (2004)): the 3D layout resembles the wiring of magnetic-core memories of the 1960s, with their bit planes corresponding to the flat dendritic trees of Purkinje cells. Kanerva (1993) illustrates step by step the cerebellum’s likeness to computer memory and to ANNs. The memory circuit consists of mossy fibers, granule cells, Purkinje cells, and climbing fibers. The axons of granule cells bifurcate into parallel fibers that run perpendicular to Purkinje cells’ flat dendritic trees.
The rest of the paper and the target article are about applying the architecture to autonomous learning by robots, modeled after learning by humans and animals with sensorimotor systems controlled by advanced brains. The paper represents a systems approach that is aligned with James Albus’s ideas about robots with human-like faculties. His book on Brains, Behavior, and Robotics (Albus 1981) came out a decade before ideas about computing with vectors in superposition began to take hold, but is still relevant and worth reading.
@@ 3 Traditional Computing with Numbers
This section serves readers who know computers from having programmed them in high-level languages but know less about the hardware within. The lesson is that the computer in itself is exceedingly simple and that its success comes from programming.
The traditional computing architecture, also known as the von Neumann architecture, consists of a memory for storing programs and data (Random-Access Memory, RAM), and a processor that steps through the program and does the math (Central Processing Unit, CPU). The memory is made of sequentially numbered storage locations, and a location’s number serves as its address. Even small computers have memories with millions of storage locations.
A storage location is a piece of hardware that can hold a small amount of data, for example a number, a letter, or a simple command for the processor to carry out, such as adding two numbers. A storage location can be visualized as a fixed number of bits, 32 for example, and the programmer decides what the bits in any particular location—at any particular address—mean.
The CPU has two functions, step-by-step running of the program and the math that is built into the computer’s circuits. The math usually includes addition, subtraction, multiplication, and division of numbers, and logic operations on strings of bits (AND, OR, NOT, shift and rotate). More complicated operations are made with programs that combine these basic operations. For example, a program to calculate the mean of a set of numbers copies the numbers, one at a time, from the memory into the processor, adds it to the sum, divides the final sum by the count, and stores the result in memory for later use—for comparing the mean heights of two groups of people, for example.
3.1 Random-Access Memory—RAM
Random Access is an all-important property of the memory. It means that the contents of any storage location are immediately available to the processor. The processor merely needs to “know” the location’s address. Random access is in contrast to storing data on tape or disk that make the processor wait for the addressed memory location to be under the read/write head. The “miracle” of random access is achieved with an address-decoder circuit.
3.2 Linked Lists
Random access allows lists of numbers to be stored so that any number in the list can start the retrieval of the rest of the list. The numbers in the list are used as addresses of storage locations and as data to store, and the list is stored by stepping through it one number at a time, and using it as the address in which to store the next number. Linked lists (also called pointer chains) are standard fare in list-processing languages such as Lisp, and provide a way to retrieve a list starting with any number in it. When vectors are linked in a list like that, they can become a predictive model of the world, as discussed under “predicting” in subsection 6.3.
@@ 4 Computing with Vectors
By a $D$-dimensional vector we mean a series of $D$ numbers treated as a coherent unit. The traditional (von Neumann) architecture for computing with numbers is readily extended to computing with wide vectors (e.g., $D$ = 10,000). The “new” architecture is essentially the old: central processing unit (CPU) to carry out the vector math, Random-Access Memory (RAM) for wide vectors, and circuits for input and output. A traditional computer would control the running of the program, with a vector-arithmetic unit running as a coprocessor, for the time being at least.
4.1 Experimenting with Vectors
We wrote a program for identifying languages using four operations on vectors: addition, multiplication, permutation, and dot product (Joshi et al. 2017). The experiment was motivated by people’s ability to identify languages by how they sound, without knowing any of the words. Rather than speech, we used about a million bytes of text for each of 21 European Union languages, transcribed in Latin alphabet of 26 letters and a symbol for the space between words. The challenge was to identify languages by how they look in print, without knowing any of the words. We made a 10,000-dimensional (10K) profile vector for each language and for each test sentence, and identified a test sentence with the language that had the most similar profile.
An analysis of the language profiles clustered them in four families: Slavic, Baltic, Romance and Germanic, with English between Romance and Germanic. Of 21,000 test sentences, 97.3% were identified correctly, and mistaken ones fell mostly in the same language family. The experiment was ran on a laptop computer in a single pass over the data and the test sentences, and took less than 8 minutes.
Profiles for languages and test sentences were computed with the same algorithm, and were compared to each other with the cosine. We started with 10K vectors for letters, computed 10K vectors for trigrams, and added them into 10K vectors for profiles. Each of the 27 letters was assigned a 10K independent random seed vector of 1s and 1-1s. The trigrams of the data are all its three-letter sequences and they were encoded into 10K trigram vectors with permutation and multiplication: permute the first letter (its 10K code) twice, permute the second letter once, take the third as is, and multiply the three. Permutation was done by rotating vector coordinates and multiplication was done coordinate by coordinate. Finally, a 10K profile vector was computed by summing over all the trigram vectors of the text: the profile and test vectors were sums of products of permutations of seed vectors for letters. The meanings of “add,” “multiply,” “permute,” and “similar” are explained next.
@@ 4.2 Instruction Set for Vectors
Three simple operations make a surprisingly powerful system to compute with: vector addition, coordinatewise multiplication, and the shuffling of vector coordinates (permutation). The vectors can be binary or integer or real or complex—the computing power comes more from high dimensionality (e.g., $D$ = 10,000) and randomness than from the nature of vector components.
The operations have been discussed at length in the literature (Plate 2003; Kanerva 2009, 2023) and can be illustrated with binary vectors (equivalent to vectors of 1s and 1-1s). Addition is by coordinatewise majority rule with ties broken at random (equivalently, ordinary vector addition followed by applying a threshold to each coordinate); multiplication is coordinatewise Exclusive-Or (XOR); and permutations reorder vector coordinates. The algebra of addition and multiplication resembles ordinary arithmetic in regard to associativity, commutativity, invertibility, and distributivity. Permutations have no counterpart in ordinary arithmetic. They are invertible, most do not commute, and they distribute over both vector addition and multiplication. Permutations of coordinates add all finite groups up to size $D$ into the vector math.
The operations allow traditional data structures (sets, sequences, lists, trees, stacks, etc.) to be represented and operated on in superposed vectors (Kleyko et al. 2022). For example, sets can be encoded with addition, the binding of variables to values with multiplication, sequences with permutations, and data structures at large with combinations of the three operations. For a simple example, a vector that means $x=a$ can be encoded by the product P=XAP=X*A where $X$ and $A$ are $D$-dimensional vectors representing the variable $x$ and the value $a$, and * is coordinatewise multiplication (XOR for binary vectors). The three operations are also used to decode a composed vector into its constituent vectors: the value of $x$ in $P$ can be recovered by multiplying $P$ with (the inverse of) $X$, namely XP=X(XA)=($X$)A=AX*P=X*(X*A)=(X*X)*A=A, and by noting that XOR is its own inverse so that XXX*X evaluates to a vector of 0s. In ordinary computing, the variable $x$ is represented by an address of a memory location and is buried in the program code, and the value $a$ is represented by the contents of the location. The three operations also allow regularities in data to be captured in vectors, opening a path to statistical learning.
@@ 4.3 Computing in Superposition
Computing in superposition goes by many names, depending mainly on the intended audience—whether mathematicians, physicist, cognitive scientists, linguists, engineers, or computer scientists: Holographic Reduced Representation (HRR), Vector Symbolic Architecture (VSA), Context-Dependent Thinning, Hyperdimensional Computing, and Semantic Pointer Architecture (SPA). Each item or element or feature or thing or piece of data—each vector—that is included in a superposed vector is distributed evenly over all vector components—any subset of vector components represents the same thing as the entire vector, only less precisely. Representation of this kind is called “holographic” or “holistic,” as opposed to the conventional localist representation where a variable is represented by an address of a memory location and its value by the contents of the location. The encoding of $x=a$ with the vector $P$ in subsection 4.2 is an example of holistic representation: P=XAP=X*A.
@@
Computing in superposition begins with $D$-dimensional seed vectors chosen at random (with i.i.d. components), to represent basic entities such as variables and names. In essence, they are symbols or tokens to compute with. They can be used as addresses to memory and as data stored in memory, and as inputs to the math operations that combine them into new vectors, which in turn can address the memory, be stored as data, and serve as inputs to the vector operations. New representations are made from existing ones with explicit calculation, which is fundamentally different from the generation of representations in an autoencoder or in the layers of a deep neural net as it is trained. Distributivity and invertibility of operations also mean that codes for composite objects can be decoded into their constituents, in some cases at least.
@@
The operations allow a high degree of parallelism. Because addition and multiplication happen coordinatewise, independently of other coordinates, all $D$ (= 10,000) can be done at the same time. Furthermore, because the operations are simple, it is feasible for each dimension to have its own processor circuit. Computing with vectors in superposition can then be seen as each coordinate having its own simple computer, all computing the same thing but at exceedingly low precision—down to 1-bit. On the aggregate, however, they compute with high precision if the dimensionality is high enough—and even if some of the simple circuits malfunction! In contrast, traditional circuits for computing with numbers are complicated and are expected to work flawlessly. The simplicity of the coordinatewise operations makes it possible to distribute them throughout the system, reducing the need to move data before it can be computed on. This is called “in-memory computing.”
5 Ideas for Architecture from Biology and Psychology
@@
How does computing with vectors fare in the light of biology and behavioral sciences? Psychological research in the 1960s and ’70s identified two levels of organization of human memory, short-term working memory and long-term data store (Atkinson & Shiffrin 1968; Anderson & Bower 1973); see also Wikipedia article on “Atkinson–Shiffrin memory model”). Representations are assembled and analyzed in the working memory and stored in the long-term memory (Tulving & Donaldson 1972; Tulving & Thomson 1973). These early models summarize behavior without accounting for underlying mechanisms and their math.
@@
The psychologists’ models are similar to traditional computing that assembles and interprets data structures in the central processing unit (CPU) and stores them in the random-access memory (RAM). Computing with vectors follows the same organization: vectors are encoded and decoded in the working memory (counterpart of CPU) and stored indefinitely in the long-term memory (counterpart of RAM). We will discuss the long-term memory first.
@@ 5.1 Long-Term Memory for Vectors
Memory allows past experience to bear on the present situation. Associative memory, also called content-addressable (these terms lack a precise definition), has been an object of psychological inquiry for over a century (James 1890; Raaijmakers & Shiffrin 1981), and mathematical models have been proposed since the early days of artificial neural nets (see Hinton & Anderson 1981, for a review). Sparse Distributed Memory (Kanerva 1988b) takes the idea a step further by assuming that the memory could have billions of locations and that its function is long-term data storage. In the rest of this paper, “long-term memory,” “associative memory,” “high-dimensional RAM,” “Sparse Distributed Memory,” and “SDM” mean the same thing and are often referred to simply as the memory. Its similarity to computer memory is significant: the computer RAM stores numbers and is addressed by numbers, an associative memory such as the SDM stores vectors and is addressed by (cued with) vectors.
Psychologically realistic models of memory for vectors, such as the Sparse Distributed Memory (Kanerva 1988b, 1993), consists of locations for storing $D$-dimensional vectors. Each location is associated with a $D$-dimensional address. A location is activated when the memory is probed/cued/addressed with a vector that is similar to the location’s address. The similarity criterion is chosen so that multiple locations are activated, but only a tiny fraction of all possible locations (Kanerva 1993). An activated location either accepts a $D$-dimensional vector (write operation) or outputs a $D$-dimensional vector (read operation). The final output of a memory read is computed from the outputs of all activated locations. Although the number of possible locations borders the incalculable, it is possible to simulate random access by distributing data over multiple locations and reconstructing it upon retrieval.
@@
There is a difference of lesser importance to note: storing a number in a computer memory replaces the old contents of the activated/addressed location; storing a vector in an associative memory adds it to the contents of all activated locations—the newly stored data are superposed.
@@
An associative memory’s likeness to the cortex of the cerebellum is remarkable (Marr 1969; Albus 1971; Kanerva 1993), and the numbers agree with the need for a high-capacity memory. With 50 billion granule cells (Llinás et al. 2004) representing locations for storing of vectors, the human brain could store one vector every minute (more than 10 million bits a day) over a life span of a 100 years.
@@
Catastrophic forgetting is a known phenomenon of artificial neural nets. It is in part due to the limit on the amount of information that any given “hardware” can store, and in part to training a monolithic network with error back-propagation (French 1999). Catastrophic means that when the capacity limit is reached, the net begins to produce outputs that are unrelated to the vectors it was trained on—it’s like a phase transition in physics where the properties of a substance change suddenly, as when super-cooled water turns into a block of ice. The alternative is graceful degradation, which can be achieved with a memory that is organized in storage locations whose contents can change but addresses are fixed, as they are in ordinary computer memory, in the Sparse Distributed Memory, and apparently also in the cortex of the cerebellum. Parallel-fiber synapses with Purkinje-cell dendrites are modifiable and make up over half the brain’s synapses; (Llinás et al. 2004). However, a small fraction of synapses (1%) should be fixed according to the theory, namely, synapses between mossy fibers and granule cells that determine which locations/granule cells to activate. They act as an address decoder and provide a framework for recording a lifetime of experience.